# FOURIER-MODULATED IMPLICIT NEURAL REPRESENTATION FOR MULTISPECTRAL SATELLITE IMAGE COMPRESSION

### v1

@@ I Introduction
In this paper, we introduce Implicit Neural Representations for Multispectral Satellite Images (ImpliSat), a novel INR-based compression framework specifically designed for MSI. ImpliSat models MSI data as coordinate-based functions, compressing and reconstructing the data while accounting for the differences in spatial resolution and channel information. Additionaly, our ImpliSat framework introduces a hypernetwork-based Fourier modulation technique that dynamically generates appropriate Fourier bases tailored to each spectral band resolution. By conditioning the modulation on the unique resolution and channel information of each band, our approach ensures that each spectral channel is represented optimally, enhancing the efficiency of compression and the accuracy of reconstruction.
@@ II Preliminary: Multispectral Satellite Images and Compression
Unlike standard RGB images, MSI encompasses multiple spectral bands, often exceeding ten. Each spectral band has different pixel value range and covers a specific range of wavelengths with different GSD depending on the sensor (cf. Fig. 1). GSD denotes the actual ground distance between adjacent pixels (10m GSD means 1 pixel in the image represents 1m in the actual distance). The bigger the GSD, the lower the spatial resolution, which means there will be fewer details visible in the image.
@@ III Proposed Methods
ImpliSat is built based on modulated INR frameworks. Our key contribution is Fourier modulation, a novel modulation technique that leverages spectral information and low-rank adaptation [17, 18, 19] to efficiently handle the multi-bands and multi-resolution nature of MSI data. The overall architecture is shown in Fig. 2. ImpliSat consists of two main components: i) Hypernetwork that generates Fourier modulations based on resolution and channel information, and ii) SIREN-based backbone INR model [8] that uses these modulation vectors to represent each band of the multispectral image.
@@ IV-B Comparison with Existing Modulated INRs
Fig. 3 presents a visualization comparison for each MSI [^2] Due to space constraints, we only show one band for each MSI in Fig. 3, where the row (top to bottom) correspond to Cairo, Merapi, London, Seoul, and Hawaii. The first and second rows show bands with 10m GSD, the third and fourth rows show bands with 20m GSD, and the last one is a band with 60m GSD. These results further illustrate that our Fourier modulation captures sharper and more detailed structures in complex areas, such as the bridges and buildings visible in the Seoul B8A and segmentation between land patches in London B7. Unlike shift and scale modulation, which tend to blur fine details, our approach preserves clear edges and sharp features, leading to significantly better reconstructions regardless of the GSD. In Fig. 4, we show the PSNR comparison between Fourier modulation and the baseline modulations throughout the training process on the Cairo dataset. Fourier modulation has much better convergence rate, as evidenced by the rapid increase in PSNR, and continues to maintain superior performance as training progresses.

### v2

@@ I Introduction
In this paper, we introduce Implicit Neural Representations for Multispectral Satellite Images (ImpliSat), a novel INR-based compression framework specifically designed for MSI. ImpliSat models MSI data as coordinate-based functions, compressing and reconstructing the data while accounting for the differences in spatial resolution and channel information. Additionally, our ImpliSat framework introduces a hypernetwork-based Fourier modulation technique that dynamically generates appropriate Fourier bases tailored to each spectral band resolution. By conditioning the modulation on the unique resolution and channel information of each band, our approach ensures that each spectral channel is represented optimally, enhancing the efficiency of compression and the accuracy of reconstruction.
@@ II Preliminary: Multispectral Satellite Images and Compression
Unlike standard RGB images, MSI encompasses multiple spectral bands, often exceeding ten. Each spectral band has different pixel value range and covers a specific range of wavelengths with different GSD depending on the sensor (cf. Fig. 1). GSD denotes the actual ground distance between adjacent pixels (10m GSD means 1 pixel in the image represents 10m in the actual distance). The bigger the GSD, the lower the spatial resolution, which means there will be fewer details visible in the image.
@@ III Proposed Methods
ImpliSat is built based on modulated INR frameworks. Our key contribution is Fourier modulation, a novel modulation technique that leverages spectral information and low-rank adaptation [17, 18, 19] to efficiently handle the multi-band and multi-resolution nature of MSI data. The overall architecture is shown in Fig. 2. ImpliSat consists of two main components: i) Hypernetwork that generates Fourier modulations based on resolution and channel information, and ii) SIREN-based backbone INR model [8] that uses these modulation vectors to represent each band of the multispectral image.
@@ IV-B Comparison with Existing Modulated INRs
Fig. 3 presents a visualization comparison for each MSI [^2] Due to space constraints, we only show one band for each MSI in Fig. 3, where the rows (top to bottom) correspond to Cairo, Merapi, London, Seoul, and Hawaii. The first and second rows show bands with 10m GSD, the third and fourth rows show bands with 20m GSD, and the last one is a band with 60m GSD. These results further illustrate that our Fourier modulation captures sharper and more detailed structures in complex areas, such as the bridges and buildings visible in the Seoul B8A and segmentation between land patches in London B7. Unlike shift and scale modulation, which tend to blur fine details, our approach preserves clear edges and sharp features, leading to significantly better reconstructions regardless of the GSD. In Fig. 4, we show the PSNR comparison between Fourier modulation and the baseline modulations throughout the training process on the Cairo dataset. Fourier modulation has much better convergence rate, as evidenced by the rapid increase in PSNR, and continues to maintain superior performance as training progresses.